OpenStudy (anonymous):

Can Anyone Help Me Check My Answers. Anyone Good In geometry because I am hopeless Please check them I will award fan and medal

OpenStudy (anonymous):

I can try :)

OpenStudy (anonymous):

x-35+x+83=180 2x+48=180 2x=132 x=66 Then plug in 66 for each of the angles: x-35 66-35=? x+83 66-83=?

OpenStudy (anonymous):

x1 = 31 x2 = - 23

OpenStudy (anonymous):

66+88 sry I typed it wrong the first time also 31 is correct for the measure of angle 1

OpenStudy (anonymous):

It's okay Haha :D 149

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

anymore questions?

OpenStudy (anonymous):

I need help with others if that's okay

OpenStudy (anonymous):

that's fine :)

OpenStudy (anonymous):

Sorry it's blue

OpenStudy (anonymous):

I don't think I have learned this yet :(

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

you would go with the opposite of what the original statement says

OpenStudy (anonymous):

ohhh , ok so c right because D seems wrong

OpenStudy (anonymous):

The complete opposite, if the team is away...

OpenStudy (anonymous):

the team will win